1 Locally linearly dependent operators

By Peter Šemrl.

Let U and V be vector spaces. Linear operators T_{1}, …, T_{n} : U → V are locally linearly dependent if for every u ∈ U the vectors T_{1}u, …, T_{n}u are linearly dependent. Some recent results on such operators will be presented.